### v1

@@ 4 Lower bound in one dimension
In the Orthogonal Vectors problem (OV) we are given two sets of vectors $X,Y\subseteq\{0,1\}^{d}$ of size $|X|=|Y|=n$ and the task is to decide whether there exist $x\in X$ and $y\in Y$ with $x\cdot y=0$, where $x\cdot y=\sum_{i=1}^{d}x[i]\cdot y[i]$. A naive algorithm solves this problem in time $\mathcal{O}(n^{2}d)$.
Hypothesis 4.1 (Orthogonal Vectors Hypothesis (OVH) [45, 44]).
@@
In this section we prove the following theorem.
@@
Assuming OVH, for any constant $\delta>0$ there is no algorithm that, given sets $B,R\subseteq\mathbb{R}$ of size $n=|R|\geq|B|=\Omega(n)$, computes $\mathrm{EMDuT}(B,R)$ in time $\mathcal{O}(n^{2-\delta})$. This even holds with the additional restriction $B,R\subseteq\{0,1,\ldots,\mathcal{O}(n^{4})\}$.
Observe that this immediately implies Theorem 1.2 because each coordinate is bounded by a polynomial. Hence, from now on, we focus on the proof of Theorem 4.2. In Section 4.1 we first construct vector gadgets. In Section 4.2 we show how, given sets $X$ and $Y$ of vectors, we use these gadgets to construct the sets of points $R$ and $B$. We prove some useful properties of these sets in Section 4.3, and finally in Section 4.4 we prove correctness of the reduction.
@@ 4.1 Vector gadgets
We have two different types of gadgets depending on whether a vector belongs to set $X$ or $Y$ (see Figure 4.1 for illustration):
@@
For a vector $x\in\{0,1\}^{d}$, create a group of points $R(x)$ to consist of: $8d$ points at the coordinate 00, and $8d$ points at the coordinate $14d+$ 1. Next, for every $i\in\{1,\ldots,d\}$: (i) if $x[i]=0$, we put points $\{4i-3,4i-2,4i-1,4i\}$, and (ii) if $x[i]=1$, we put points $\{4i-2,4i-1\}$.
@@
For a vector $y\in\{0,1\}^{d}$, create a group of points $B(y)$: (i) if $y[i]=0$ put points $\{4i-2,4i-1\}$, and (ii) if $y[i]=1$ put points $\{4i-3,4i\}$.
Figure 4.1: Gadgets for red and blue vectors in $d=2$. The top figure shows $R(x)$ for $x=(1,0)$, and the bottom figure illustrates $B(y)$ for $y=(0,1)$. Since $x$ and $y$ are orthogonal, each blue point corresponds to a red point with the same coordinate.
@@
Let $x,y\in\{0,1\}^{d}$ be $d$-dimensional vectors.
@@
    If $x$ and $y$ are not orthogonal then $\mathrm{EMD}(B(y)+\tau,R(x))\geq 1$ for all $\tau\in\mathbb{R}$.
Moreover, for every $\tau\in\mathbb{R}$, we have $\mathrm{EMD}(B(y)+\tau,R(x))\geq|\tau|$. If $|\tau|\geq 4d+1$, then we even have $\mathrm{EMD}(B(y)+\tau,R(x))=|\tau|\cdot c_{1}+c_{2}$, where $c_{1}=2d$ and $c_{2}=-4d^{2}-d$.
@@
For the proof of Property 1, assume that $x$ and $y$ are orthogonal. We construct a matching such that $\mathrm{EMD}(B(y),R(x))=0$. Fix an index $i\in\{1,\ldots,d\}$. If $y[i]=1$, then $x[i]=0$; therefore, points at the coordinates $\{4i-3,4i\}$ exist in $R(x)$ and we can match points of $B(y)$ at these coordinates. If $y[i]=0$, then notice the points at coordinates $\{4i-2,4i-1\}$ of $R(x)$ always exist. Hence, we can match points of $B(y)$ at these coordinates precisely.
Next, we prove Property 2. Consider any $\tau\in\mathbb{R}$. If $|\tau|>1$, it is addressed by the last property, so assume $|\tau|\leq 1$. One of the leftmost or rightmost points of $B(y)$ must be matched by an edge of length $|\tau|$ (in the case when $\tau=0$, all of them are matched exactly). Assume that $x$ and $y$ are not orthogonal. This means that there exists an $i\in\{1,\ldots,d\}$ such that $x[i]=y[i]=1$. By construction, $B(y)$ contains a point at the coordinate $4i$, but $R(x)$ does not contain such a point. Because the smallest distance between points is $|1-\tau|$, it follows that $\mathrm{EMD}(B(y),R(x))\geq|1-\tau|+|\tau|\geq 1$.
Finally, we focus on the proof of the last property. Consider any $\tau\in\mathbb{R}$. If $\tau\neq 0$, then either the leftmost or rightmost $2d$ points of $B(y)$ are not matched to the point on the same coordinate as them. Therefore, $\mathrm{EMD}(B(y)+\tau,R(x))\geq|\tau|$. Now, consider $\tau$ with $|\tau|\geq 4d+1$. Observe that $B(y)+\tau$ contains $2d$ points in total. The closest point in $R(x)$ to each of them is at the coordinate 00 or $14d+$ 1. Because there are $8d$ such points, there are enough points to match each point in $B(y)$ to the closest point in $R(x)$. Hence, $\mathrm{EMD}(B(y)+\tau,R(x))$ matches each point of $B(y)$ to a point at the coordinate 00 or $14d+$ 1 of $R(x)$. If $\tau>4d+1$, the cost of the matching is
$\mathrm{EMD}(B(y)+\tau,R(x))=\sum_{i=1}^{d}\tau-(4d+1)+(4i-2-y[i])+\sum_{i=1}^{d}\tau-(4d+1)+(4i-1+y[i]).$
This is equal to $\mathrm{EMD}(B(y)+\tau,R(x))=2d\tau-4d^{2}-d$. Consequently, we can take $c_{1}=2d$ and $c_{2}=-4d^{2}-d$. It remains to check whether the above choice is correct for $\tau<-(4d+1)$. In that case, the cost of the matching is
$\mathrm{EMD}(B(y)+\tau,R(x))=\sum_{i=1}^{d}|\tau|-(4i-2-y[i])+\sum_{i=1}^{d}|\tau|-(4i-1+y[i]).$
This, again, after simplification is equal to $2d|\tau|-4d^{2}-d$ and the same choice of $c_{1}$ and $c_{2}$ works. ∎
@@ 4.2 Reduction
Now we use the vector gadgets from the previous section to reduce from the Orthogonal Vectors problem to $\mathrm{EMDuT}$. Specifically, given an OV instance $X,Y\subseteq\{0,1\}^{d}$ of size $n$, we construct sets $B,R\subseteq\mathbb{R}$ such that from $\mathrm{EMDuT}(B,R)$ we can easily infer whether $X,Y$ contains an orthogonal pair of vectors or not. Our reduction takes time $\mathcal{O}(nd)$ to construct the sets $B,R$, in particular the constructed sets have size $\mathcal{O}(nd)$. Hence, if there would be an algorithm computing $\mathrm{EMDuT}(B,R)$ in time $\mathcal{O}(|R|^{2-\delta})$ for some constant $\delta>0$, then our reduction would yield an algorithm for OV running in time $\mathcal{O}((nd)^{2-\delta})$, which contradicts OVH (Hypothesis 4.1). That is, assuming OVH, $\mathrm{EMDuT}(B,R)$ cannot be computed in time $\mathcal{O}(|R|^{2-\delta})$ for any constant $\delta>0$.
For the reduction we can assume that $n$ is odd, because if $n$ is even, one can simply add a vector consisting exclusively of $1$s to both $X$ and $Y$. We can also assume that $d\leq n$, since otherwise the naive algorithm for OV already runs in time $\mathcal{O}(n^{2}d)=\mathcal{O}(nd^{2})$. Our reduction constructs the following point sets, for $\Delta\coloneqq 1000dn$:
  • Red Points: For the $i^{\mathrm{th}}$ vector $x_{i}\in X$, we create five red gadgets $R(x_{i})^{(1)},\ldots,R(x_{i})^{(5)}$. For each $k\in[5]$, we translate $R(x_{i})^{(k)}$ by $(i+kn)\cdot(n-1)\Delta$ and call it $(i+kn)^{\mathrm{th}}$ red cell.
@@
We create five copies of red points for a technical reason that will become clear later (just three copies is enough, but then we would need to argue about two types of optimal translations in the analysis). We denote the set of all red points by $R$, and the set of all blue points by $B$. This concludes the construction. Observe that $B,R$ can be constructed in time $\mathcal{O}(nd)$, as claimed, and that their coordinates are in $\{0,\ldots,\mathcal{O}(dn^{3})\}\subseteq\{0,\ldots,\mathcal{O}(n^{4})\}$. Let $c_{1}$ and $c_{2}$, where $c_{1},c_{2}$ are the constants (that depend on $d$) from Lemma 4.5. Let
$\Lambda\coloneqq c_{1}\Delta\cdot(n^{2}-1)/4+c_{2}\cdot(n-1).$
@@ 4.3 Properties of the construction
We now prove several useful properties about the construction described above.
@@
There exist $i_{\tau},j_{\tau}\in\{0,\ldots,n-1\}$ and $\varepsilon\in(-1/2,1/2]$ such that:
@@
such that $\mathrm{EMD}(B+\tau,R)=\mathrm{EMD}(B,R)$.
@@
Let $\lambda\coloneqq(n-1)n\Delta$. We will show that there exists an $\tau\in[\lambda,3\lambda]$ such that $\mathrm{EMD}(B+\tau,R)=\mathrm{EMD}(B,R)$. This will conclude the proof, because (i) (4.1) is minimized for $i_{\tau}=0$ and $j_{\tau}=n-1$ (ii) is maximized when $j_{\tau}=0$ and $i_{\tau}=n-1$, and (iii) steps of $i_{\tau},j_{\tau}$ are $\Delta$ and $\varepsilon\Delta\in[-\Delta/2,\Delta/2]$. Hence every real number in $[\lambda,3\lambda]$ is covered by some choice of $i_{\tau},j_{\tau}$ and $\varepsilon$.
If $\tau<\lambda$, then we claim that $\mathrm{EMD}(B+\tau,R)\geq\mathrm{EMD}(B+\tau+\lambda,R)$. This holds, because (i) the right endpoint of $B+\tau+\lambda$ is on the left of right endpoint of $R$, and (ii) because $R$ is periodic, with period $\lambda$, each edge of the matching $\mathrm{EMD}(B+\tau,R)$ can be shifted by $\lambda$.
An analogous argument ensures that when $\tau>3\lambda$, then $\mathrm{EMD}(B+\tau,R)\geq\mathrm{EMD}(B+\tau-\lambda,R)$. ∎
From now, we assume that our optimal translation is of the form (4.1) and the parameters $i_{\tau},j_{\tau}$ and $\varepsilon$ are known. The idea behind this form is to have the property that the $i_{\tau}^{\mathrm{th}}$ red cell and $j_{\tau}^{\mathrm{th}}$ blue cell “nearly-align”. Before we explain it in more details, our goal is to show that $|\varepsilon|$ is small.
@@
If $j_{\tau}\in\{0,\ldots,n-1\}$, then for every $\varepsilon\in(-1/2,1/2]$ it holds that:
$\sum_{k=0}^{n-1}\min\{\lvert k-j_{\tau}+\varepsilon\rvert,\lvert n-1+k-j_{\tau}+\varepsilon\rvert,\lvert n-1-k+j_{\tau}-\varepsilon\rvert\}=\lvert\varepsilon\rvert+(n^{2}-1)/4.$
@@
First, consider $j\leq(n-1)/2$. Then $\lvert k-j\rvert\leq\lvert n-1+j-j\rvert$ and we need to show:
$\sum_{k=0}^{n-1}\min\{\lvert k-j+\varepsilon\rvert,\lvert n-1-k+j-\varepsilon\rvert\}=\lvert\varepsilon\rvert+(n^{2}-1)/4.$ (4.2)
The proof when $j\geq(n+1)/2$ is analogous, as then $\lvert k-j\rvert\geq\lvert n-1+k-j\rvert$. Then, one can take $j^{\prime}=j-(n-1)/2$, observe that $\lvert n-1+k-j\rvert=\lvert k-j^{\prime}\rvert$ and repeat our arguments. Hence, we focus on proving (4.2) under the assumption that $j\leq(n-1)/2$. The sum in (4.2) sum can be stratified into:
$|\varepsilon|+\sum_{k=0}^{j-1}\lvert k-j+\varepsilon\rvert+\sum_{k=j+1}^{(n-1)/2+j}\lvert k-j+\varepsilon\rvert+\sum_{k=(n+1)/2+j}^{n}\lvert n-1-k+j-\varepsilon\rvert.$
Now, we can deduce the sign of each absolute value, so:
$|\varepsilon|+\sum_{k=0}^{j-1}(j-k-\varepsilon)+\sum_{k=j+1}^{(n-1)/2+j}(k-j+\varepsilon)+\sum_{k=(n+1)/2+j}^{n}(n-1-k+j-\varepsilon).$
Observe, that the number of $\varepsilon$ with positive and negative signs is equal $(n-1)/2$, hence these cancel out and:
$|\varepsilon|+\sum_{k=0}^{j-1}(j-k)+\sum_{k=j+1}^{(n-1)/2+j}(k-j)+\sum_{k=(n+1)/2+j}^{n}(n-1-k+j).$
@@
$|\varepsilon|+\sum_{\ell=1}^{j}\ell+\sum_{\ell=1}^{(n-1)/2}\ell+\sum_{\ell=j+1}^{(n-1)/2}\ell=|\varepsilon|+2\sum_{\ell=1}^{(n-1)/2}\ell=|\varepsilon|+(n^{2}-1)/4.$
Here, we used the formula $\sum_{i=1}^{m}i=m(m+1)/2$ for $m=(n-1)/2$. ∎
@@
$\min_{\ell\in\mathbb{Z}}\left\{\lvert k\cdot n\Delta+\tau-\ell\cdot(n-1)\Delta\rvert-2(4d+1)\right\},$
because the width of every cell is $14d+$ 1. Note that the number of points in each blue cell is exactly $2d$, hence:
$\textup{LHS}\coloneqq\mathrm{EMD}(B+\tau,R)$ $\geq\sum_{k=0}^{n-1}2d\cdot\left(\min_{\ell\in\mathbb{Z}}\left\{\lvert kn\cdot\Delta+\tau-\ell(n-1)\cdot\Delta\rvert\right\}-2(4d+1)\right).$
Next, we plug in the definition of $\tau$:
LHS $\geq-4d(4d+1)+2d\Delta\cdot\sum_{k=0}^{n-1}\min_{\ell\in\mathbb{Z}}\left\{\lvert(n-1)(2n+i_{\tau}-j_{\tau}+k-\ell)+k-j_{\tau}+\varepsilon\rvert\right\}.$
@@
LHS $\geq-4d(4d+1)+2d\Delta\cdot\sum_{k=0}^{n-1}\min\left\{\lvert k-j_{\tau}+\varepsilon\rvert,\lvert n-1+k-j_{\tau}+\varepsilon\rvert,\lvert n-1-k+j_{\tau}-\varepsilon\rvert\right\}.$
Next, we recall that $j\in\{0,\ldots,n-1\}$ and use Proposition 4.7:
LHS $\geq-4d(4d+1)+2d\Delta\cdot\left(|\varepsilon|+(n^{2}-1)/4\right).$
Finally, we use $0.1\leq|\varepsilon|\leq 0.5$ and the fact that $c_{1}=2d$:
LHS $\geq-4d(4d+1)+2d\Delta\lvert\varepsilon\rvert+\Lambda-(n-1)c_{2}>\Lambda.\qed$
Consider a cell $k\in\{0,\ldots,n-1\}$ of $B$. We define $\textup{NearestCell}(k)$ as:
$\textup{NearestCell}(k)=\begin{cases}2n+i_{\tau}-j_{\tau}+k&\text{if }|k-j_{\tau}|<n/2,\\ 2n+i_{\tau}-j_{\tau}+k+1&\text{if }k-j_{\tau}>n/2,\\ 2n+i_{\tau}-j_{\tau}+k-1&\text{if }k-j_{\tau}<-n/2.\\ \end{cases}$
Recall that $n$ is odd, and hence all the cases are covered. We show that $\textup{NearestCell}(k)$ is the index of the red cell that is closest to the $k^{\mathrm{th}}$ cell of $B$ after translation $\tau$, meaning that the optimal matching matches all points in the $k^{\mathrm{th}}$ blue cell to points in the $\textup{NearestCell}(k)^{\mathrm{th}}$ red cell:

### v2

@@ 4 Lower Bound in One Dimension
In the Orthogonal Vectors problem (OV), we are given two sets of vectors $X,Y\subseteq\{0,1\}^{d}$ with $n=|X|=|Y|$ and the task is to decide whether there exist $x\in X$ and $y\in Y$ with $x\cdot y=0$, where $x\cdot y=\sum_{i=1}^{d}x[i]\cdot y[i]$. A naive algorithm solves this problem in time $\mathcal{O}(|X|^{2}d)$.
Hypothesis 4.1 (Orthogonal Vectors Hypothesis (OVH) [ov-seth, icm-survey]).
@@
In this section, we prove the following theorem.
@@
Assuming OVH, for any constant $\delta>0$ there is no algorithm that, given sets $B,R\subseteq\mathbb{R}$ of size $|R|\geq|B|=\Omega(n)$, computes $\mathrm{EMDuT}(B,R)$ in time $\mathcal{O}(n^{2-\delta})$. This even holds with the additional restriction $B,R\subseteq\{0,1,\ldots,\mathcal{O}(n^{4})\}$.
Observe that this immediately implies Theorem˜1.2 because each coordinate is bounded by a polynomial. Hence, from now on, we focus on the proof of Theorem˜4.2.
We start by briefly sketching the reduction. As a building block, we design vector gadgets that for two vectors $x\in X\subseteq\{0,1\}^{d}$ and $y\in Y\subseteq\{0,1\}^{d}$ are sets of points $B(x)$ and $R(y)$ such that $\mathrm{EMDuT}(B(x),R(y))=0$ if $x$ and $y$ are orthogonal and $\mathrm{EMDuT}(B(x),R(y))\geq 1$ otherwise. This gadget is constructed by encoding $x[i]$ and $y[i]$ coordinate-wise with the geometric patterns presented in Figure˜4.1. Importantly, these gadgets share the same barycentre, therefore, when two such patterns are translated far apart (i.e., their convex hulls do not intersect), then their $\mathrm{EMD}$ does not depend on the coordinates of the vectors. In Section˜4.1 we formally construct these vector gadgets.
This allows us to construct a gadget such that for each translation $\tau$ only a single pair of vector gadgets overlaps, and the $\mathrm{EMD}$ between any other pair of vector gadgets does not depend on the coordinates of the vectors. In Section˜4.2 we show construct $B$ and $R$ as such collection of appropriately-spaced copies of gadgets $B(x)$ and $R(y)$ for $x\in X$ and $y\in Y$ in such a way that there exists a threshold $\Lambda$ such that $\mathrm{EMDuT}(B,R)=\Lambda$ if there are two orthogonal vectors in $X\times Y$ and $\mathrm{EMDuT}(B,R)\geq\Lambda+1$ otherwise. We formally prove these properties in Sections˜4.4 and 4.3.
@@ 4.1 Vector Gadgets
We construct two different types of gadgets depending on whether a vector belongs to set $X$ or $Y$ (see Figure˜4.1 for illustration):
@@
For any vector $x\in\{0,1\}^{d}$, define a set of points $R(x)$ to consist of:
  • $8d$ points at the coordinate 0,
  • $8d$ points at the coordinate $4d+1$, and
  • for every $i\in\{1,\ldots,d\}$:
    • if $x[i]=0$, add points $\{4i-3,4i-2,4i-1,4i\}$,
    • if $x[i]=1$, add points $\{4i-2,4i-1\}$.
@@
For any vector $y\in\{0,1\}^{d}$, define a set of points $B(y)$ to consist of:
  • one point at the coordinate 0,
  • one point at the coordinate $4d+1$, and
  • for every $i\in\{1,\ldots,d\}$:
    • if $y[i]=0$, add points $\{4i-2,4i-1\}$,
    • if $y[i]=1$, add points $\{4i-3,4i\}$.
Figure 4.1: Gadgets for red and blue vectors in $d=2$. The top figure shows $R(x)$ for $x=(1,0)$, and the bottom figure shows $B(y)$ for $y=(0,1)$. Since $x$ and $y$ are orthogonal, each blue point corresponds to a red point with the same coordinate.
@@
Let $w=4d+1$ be the width of the gadget. Let $x,y\in\{0,1\}^{d}$ be $d$-dimensional vectors.
@@
    If $x$ and $y$ are not orthogonal then $\mathrm{EMD}(B(y)+\tau,R(x))\geq\max\{1,|\tau|\}$ for all $\tau\in\mathbb{R}$.
Moreover, if $|\tau|\geq w$, then we have $\mathrm{EMD}(B(y)+\tau,R(x))=|\tau|\cdot c_{1}-c_{2}$, where $c_{1}=2(d+1)$ and $c_{2}=4d^{2}+7d+1$.
@@
For the proof of Property 1, assume that $x$ and $y$ are orthogonal. We construct a matching such that $\mathrm{EMD}(B(y),R(x))=0$. Fix an index $i\in\{1,\ldots,d\}$. If $y[i]=1$, then $x[i]=0$; therefore, points at the coordinates $\{4i-3,4i\}$ exist in $R(x)$ and we can match the points of $B(y)$ at these coordinates. If $y[i]=0$, then note that the points at coordinates $\{4i-2,4i-1\}$ of $R(x)$ always exist. Hence, we can precisely match the points of $B(y)$ at these coordinates.
Next, we prove Property 2. Consider any $\tau\in\mathbb{R}$. At least one point of $B(y)$ (namely the leftmost or the rightmost point) must be matched by an edge of length $|\tau|$. This proves Property 2 if $|\tau|\geq 1$. So consider the case $|\tau|<1$. If $x$ and $y$ are not orthogonal, there exists an $i\in\{1,\ldots,d\}$ such that $x[i]=y[i]=1$. By construction, $B(y)$ contains points at the coordinates $4i$ and 4i34i-3, but $R(x)$ does not contain such points. Thus, the point $4i$ in $B(y)$ is matched to a point in distance at least $1-|\tau|$. In addition to the cost of $|\tau|$ incurred by the leftmost or rightmost point of $B(y)$, we obtain $\mathrm{EMD}(B(y),R(x))\geq 1$.
Finally, we focus on the proof of the last property. Consider any $\tau\in\mathbb{R}$ with $|\tau|\geq w$. Observe that $B(y)+\tau$ contains $2(d+1)$ points in total. The closest point in $R(x)$ to each of them is at the coordinate 0 or $w$. Because there are $8d$ such points, there are enough points to match each point in $B(y)$ to the closest point in $R(x)$. Hence, $\mathrm{EMD}(B(y)+\tau,R(x))$ matches each point of $B(y)$ to a point at the coordinate 0 or $w$ of $R(x)$. If $\tau\geq w$, the cost of the matching is
$\mathrm{EMD}(B(y)+\tau,R(x))$ $=2\tau-w+\sum_{i=1}^{d}\tau-w+(4i-2-y[i])+\sum_{i=1}^{d}\tau-w+(4i-1+y[i])$
$=2(d+1)\tau-4d^{2}-7d-1.$
Consequently, we can take $c_{1}=2(d+1)$ and $c_{2}=4d^{2}+7d+1$. It remains to check that the above choice is correct for $\tau\leq-w$. In that case, the cost of the matching is
$\mathrm{EMD}(B(y)+\tau,R(x))$ $=2|\tau|-w+\sum_{i=1}^{d}|\tau|-(4i-2-y[i])+\sum_{i=1}^{d}|\tau|-(4i-1+y[i])$
$=2(d+1)|\tau|-4d^{2}-7d-1.$
Thus, the same choice of $c_{1}$ and $c_{2}$ works. ∎
@@ 4.2 Reduction
Now we use the vector gadgets from the previous section to reduce the Orthogonal Vectors problem to $\mathrm{EMDuT}$. Specifically, given an OV instance $X,Y\subseteq\{0,1\}^{d}$ such that |X|=|Y|=n1|X|=|Y|=n-1, we construct sets $B,R\subseteq\mathbb{R}$ such that from $\mathrm{EMDuT}(B,R)$ we can easily infer whether $X,Y$ contains an orthogonal pair of vectors or not. Our reduction takes time $\mathcal{O}(nd)$ to construct the sets $B,R$, in particular the constructed sets have size $\mathcal{O}(nd)$. Hence, if there was an algorithm computing $\mathrm{EMDuT}(B,R)$ in time $\mathcal{O}(|R|^{2-\delta})$ for some constant $\delta>0$, then our reduction would yield an algorithm for OV running in time $\mathcal{O}((nd)^{2-\delta})$, which contradicts OVH (Hypothesis 4.1). That is, assuming OVH, $\mathrm{EMDuT}(B,R)$ cannot be computed in time $\mathcal{O}(|R|^{2-\delta})$ for any constant $\delta>0$.
For the reduction, we can assume that $n$ is even, because otherwise we can add a vector consisting exclusively of $1$s to both $X$ and $Y$. We can also assume that $d\leq n$, since otherwise the naive algorithm for OV already runs in time $\mathcal{O}(n^{2}d)=\mathcal{O}(nd^{2})$. Our reduction constructs the following point sets, for $\Delta\coloneqq 1000dn$:
  • Red Points: For the $i^{\mathrm{th}}$ vector $x_{i}\in X$, we create five red gadgets $R(x_{i})^{(1)},\ldots,R(x_{i})^{(5)}$. For each $k\in[5]$, we translate $R(x_{i})^{(k)}$ by $(i+kn)\cdot(n-1)\Delta$ and call it the $(i+kn)^{\mathrm{th}}$ red cell.
@@
We create five copies of red points for a technical reason that will become clear later (just three copies would be enough, but then we would need to argue about two types of optimal translations in the analysis). We denote the set of all red points by $R$, and the set of all blue points by $B$. This concludes the construction. Observe that $B,R$ can be constructed in time $\mathcal{O}(nd)$, as claimed, and that their coordinates are in $\{0,\ldots,\mathcal{O}(dn^{3})\}\subseteq\{0,\ldots,\mathcal{O}(n^{4})\}$. Let $c_{1}$ and $c_{2}$ be the constants (that depend on $d$) from Lemma˜4.5. Let
$\Lambda\coloneqq c_{1}\Delta\cdot n(n-2)/4+c_{2}\cdot(n-2).$
@@ 4.3 Properties of the Construction
We now prove several useful properties of the construction described above.
@@
There exist $i_{\tau},j_{\tau}\in\{0,\ldots,n-1\}$ and $\varepsilon\in(-1/2,1/2]$ such that for
@@
we have $\mathrm{EMD}(B+\tau,R)=\mathrm{EMDuT}(B,R)$.
@@
Let $\lambda\coloneqq(n-1)n\Delta$ and let $\tau$ be the optimal translation. If $\tau<\lambda$, then we claim that $\mathrm{EMD}(B+\tau,R)\geq\mathrm{EMD}(B+\tau+\lambda,R)$. This holds because (i) the right endpoint of $B+\tau+\lambda$ is to the left of right endpoint of $R$, and (ii) because $R$ is periodic, with period $\lambda$, each edge of the matching $\mathrm{EMD}(B+\tau,R)$ can be shifted by $\lambda$ (or matched to a strictly closer point, if the shifted point is unavailable).
An analogous argument ensures that when $\tau>2\lambda$, then $\mathrm{EMD}(B+\tau,R)\geq\mathrm{EMD}(B+\tau-\lambda,R)$. Hence we can assume that $\tau\in[\lambda,2\lambda]$. Similar arguments show also show that for $\tau\in[\lambda,2\lambda]$ we have $\mathrm{EMD}(B+\tau,R)=\mathrm{EMD}(B+\tau+\lambda,R)$.
Observe, that for $\tau\in[\lambda,2\lambda]$, there exist $k,j\in\{0,\ldots,n-1\}$ and $\varepsilon\in(-1/2,1/2]$ such that:
$\tau=((n-1)\cdot(n+k+1)-j+\varepsilon)\Delta.$
Let $i_{1}:=k+j-n+1$ and $i_{2}:=k+j+1$ and observe that either $i_{1}$ or $i_{2}$ is in $\{0,\ldots,n-1\}$ and is a valid choice of $i$. ∎
From now, we assume that our optimal translation is of the form (4.1) and the parameters $i_{\tau},j_{\tau}$ and $\varepsilon$ are known. The idea behind this form is to have the property that the $i_{\tau}^{\mathrm{th}}$ red cell and $j_{\tau}^{\mathrm{th}}$ blue cell “nearly-align”. Before we explain this in more detail, our goal is to show that $|\varepsilon|$ is small.
@@
For every even integer $n$, any $j_{\tau}\in\{0,\ldots,n-1\}$ and $\varepsilon\in(-1/2,1/2]$ it holds that:
$\sum_{k\in\{1,\ldots,n-1\}}\min\{\lvert k-j_{\tau}+\varepsilon\rvert,\lvert k-j_{\tau}+(n-1)+\varepsilon\rvert,\lvert k-j_{\tau}-(n-1)+\varepsilon\rvert\}=\lvert\varepsilon\rvert+n(n-2)/4.$
@@
We will consider the case $j_{\tau}<n/2$ as otherwise the reasoning is analogous. Then $\lvert k-j_{\tau}\rvert\leq\lvert k-j_{\tau}+n-1\rvert$ and we need to show:
$\sum_{k=1}^{n-1}\min\{\lvert k-j_{\tau}+\varepsilon\rvert,\lvert k-j_{\tau}-(n-1)+\varepsilon\rvert\}=\lvert\varepsilon\rvert+n(n-2)/4.$ (4.2)
The sum in (4.2) can be stratified into:
$|\varepsilon|+\sum_{k=1}^{j_{\tau}-1}\lvert k-j_{\tau}+\varepsilon\rvert+\sum_{k=j_{\tau}+1}^{n/2+j_{\tau}-1}\lvert k-j_{\tau}+\varepsilon\rvert+\sum_{k=n/2+j_{\tau}}^{n-1}\lvert k-j_{\tau}-(n-1)+\varepsilon\rvert$.
Now, we can deduce the sign of each absolute value, so we obtain:
$|\varepsilon|+\sum_{k=1}^{j_{\tau}-1}(j_{\tau}-k-\varepsilon)+\sum_{k=j_{\tau}+1}^{n/2+j_{\tau}-1}(k-j_{\tau}+\varepsilon)+\sum_{k=n/2+j_{\tau}}^{n-1}(n-1-k+j_{\tau}-\varepsilon).$
Observe, that the number of $\varepsilon$ with positive and negative signs is equal, hence these cancel out, which yields:
$|\varepsilon|+\sum_{k=1}^{j_{\tau}-1}(j_{\tau}-k)+\sum_{k=j_{\tau}+1}^{n/2+j_{\tau}-1}(k-j_{\tau})+\sum_{k=n/2+j_{\tau}}^{n-1}(n-1-k+j_{\tau}).$
@@
$|\varepsilon|+\sum_{\ell=1}^{j_{\tau}-1}\ell+\sum_{\ell=1}^{n/2-1}\ell+\sum_{\ell=j_{\tau}}^{n/2-1}\ell=|\varepsilon|+2\sum_{\ell=1}^{n/2-1}\ell=|\varepsilon|+n(n-2)/4.$
Here, we used the formula $\sum_{i=1}^{m}i=m(m+1)/2$ for m=n/21m=n/2-1. ∎
@@
$\min_{\ell\in\mathbb{Z}}\left\{\lvert k\cdot n\Delta+\tau-\ell\cdot(n-1)\Delta\rvert-2w\right\},$
because the width of every cell is $w=4d+1$. Note that the number of points in each blue cell is exactly $2(d+1)$, hence:
$\textup{LHS}\coloneqq\mathrm{EMD}(B+\tau,R)$ $\geq\sum_{k=1}^{n-1}2(d+1)\cdot\left(\min_{\ell\in\mathbb{Z}}\left\{\lvert kn\cdot\Delta+\tau-\ell(n-1)\cdot\Delta\rvert\right\}-2w\right).$
Next, we plug in the definition of $\tau:=((n-1)(2n+i_{\tau})-nj_{\tau}+\varepsilon)\Delta$:
LHS $\geq-4(d+1)w+2(d+1)\Delta\cdot\sum_{k=1}^{n-1}\min_{\ell\in\mathbb{Z}}\left\{\lvert(n-1)(2n+i_{\tau}-j_{\tau}+k-\ell)+k-j_{\tau}+\varepsilon\rvert\right\}$.
@@
LHS $\geq-4(d+1)w+2(d+1)\Delta\cdot\sum_{k=1}^{n-1}\min\left\{\lvert k-j_{\tau}+\varepsilon\rvert,\lvert k-j_{\tau}+n-1+\varepsilon\rvert,\lvert k-j_{\tau}-(n-1)+\varepsilon\rvert\right\}$.
Next, we use Proposition˜4.7:
LHS $\geq-4(d+1)w+2(d+1)\Delta\cdot\left(|\varepsilon|+n(n-2)/4\right).\$
Finally, we use $0.1\leq|\varepsilon|\leq 0.5$ and the fact that $\Lambda=c_{1}\Delta n(n-2)/4+c_{2}(n-2)$, where $c_{1}=2(d+1)$:
LHS $\geq-4(d+1)w+2(d+1)\Delta\lvert\varepsilon\rvert+\Lambda-(n-2)c_{2}>\Lambda,$
where we used $|\varepsilon|\geq 0.1$ and $\Delta\gg 4dn$. ∎
Consider a cell $k\in\{1,\ldots,n-1\}$ of $B$. We define $\textup{NearestCell}(k)$ as:
[…]
Recall that $n$ is even, and hence all the cases are covered. We show that $\textup{NearestCell}(k)$ is the index of the red cell that is closest to the $k^{\mathrm{th}}$ cell of $B$ after translation $\tau$, meaning that the optimal matching matches all points in the $k^{\mathrm{th}}$ blue cell to points in the $\textup{NearestCell}(k)^{\mathrm{th}}$ red cell: